A car starts from rest and accelerates for 3.2 s
with an acceleration of 2.9 m/s
2
.
How far does it travel?
Answer in units of m
1
Expert's answer
2012-09-14T12:11:40-0400
The distance traveled by car is

L = at²/2 = 2.9*3.2²/2 = 14.848 m.